OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $6.10. If the number of dimes was tripled and the number of nickels was decreased by 49, the value of the coins would be $14.05. How many nickels and dimes does he have? ----

OpenStudy (anonymous):

let x be the number of dimes and y be the nickles x+y=6.10 3x+y-49=14.02

OpenStudy (anonymous):

solve it to get the answer

OpenStudy (anonymous):

solve it for me

OpenStudy (saifoo.khan):

lol, hashir.. solve.

OpenStudy (anonymous):

awwwwww!!!!!!!!!!!!!!!11

OpenStudy (anonymous):

>.> well show them how to solve it XD

OpenStudy (saifoo.khan):

hashir baita, aap kuch bhool rahay hain

OpenStudy (anonymous):

x+y=6.10 3x+y-49=14.05

OpenStudy (anonymous):

solve it urself honey girl!!!!!!!

OpenStudy (anonymous):

yaad a gaya

OpenStudy (saifoo.khan):

baar baar na bhola karo!

OpenStudy (anonymous):

i cannot get the second equation

OpenStudy (anonymous):

3x=dimes tripled y-49=reduced by 49

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

:P for the first equation subtract x or y, I'll solve for x since it's easier IMO. x = 6.10-y now plug in your new x into the second equation 3(6.10-y)+y-49 = 14.02 can you solve from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

why not D:

OpenStudy (anonymous):

tell me what you think you should do now

OpenStudy (anonymous):

well you have to distribute 3. which means multiply 3 by everything in the ( ) so 3x6.10 and 3x-y so you get 18.30-3y+y-49=14.02

OpenStudy (anonymous):

do you know how to combine like terms?